OpenStudy (anonymous):

Evaluate: -4 - (-10)

OpenStudy (anonymous):

6

OpenStudy (aroub):

= 6

OpenStudy (anonymous):

What is the slope and y-intercept of the graph of 2x + 3y = 18?

OpenStudy (aroub):

3y=-2x+18 y=-2/3x+6 slope= -2/3 y-intercept= 6